OpenStudy (anonymous):

A bacteria population triples every 12 days. What is a possible equation for its growth?

OpenStudy (anonymous):

p1= po(x/4) :D.. just a try.. where x is the number of days

OpenStudy (amistre64):

yeah, i couldnt make sense of any of the options you provided :)

OpenStudy (anonymous):

what are the options??

OpenStudy (amistre64):

A*e^(.0196*t) if anything

OpenStudy (anonymous):

can someone tell me if my equation is right?!

OpenStudy (anonymous):

:( sorry guys. here are the options: f(t)=a*(1/3)^t/12 f(t)=a*(3)^t f(t)=a*(3)^t/12 f(t)=a*(3)^-t/12

OpenStudy (anonymous):

third option should be the answer!

OpenStudy (anonymous):

Thank you mashy! :)

OpenStudy (amistre64):

so it was a recursive equation?

OpenStudy (anonymous):

i mean first and 4th option!

OpenStudy (anonymous):

1st- f(t)=a*(1/3)^t/12 4th- f(t)=a*(3)^-t/12

OpenStudy (anonymous):

fourth is negative t/12

OpenStudy (anonymous):

So... ? Which one?

OpenStudy (anonymous):

3rd is the answer.. arre first and second one are the same (1/3)=3^-1

OpenStudy (anonymous):

K. Thanks mashy :)

OpenStudy (anonymous):

again i mean't FIRST AND FOURTH.. gosh i dunno whats wrong with me!